9/4 + 3y equals 6y - 3/20
Answer by ReadingBoosters(3246)

Starting with
9/4 + 3y = 6y - 3/20
Isolate y
9/4 + 3/20 + 3y - 3y = 6y - 3/20 + 3/20 - 3y
9/4 + 3/20 = 3y
Set common denominator of 20
9/4(5/5) + 3/20 = 3y
45/20 + 3/20 = 3y
48/20 = 3y
(48/20)/3 = 3y/3
same as (48/20)(3) = y
144/20 = y
7 4/20 = 7 1/5 = y